### v1

# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

Abstract
Code watermarking identifies AI-generated code by embedding patterns into the code during generation. Effective watermarking requires meeting two key conditions: the watermark should be reliably detectable, and the code should retain its original functionality. However, existing methods often modify tokens that are critical for program logic, such as keywords in conditional expressions or operators in arithmetic computations. These modifications can cause syntax errors or functional failures, limiting the practical use of watermarking. We present STONE, a method that preserves functional integrity by selectively inserting watermarks only into non-syntax tokens. By excluding tokens essential for code execution, STONE minimizes the risk of functional degradation.
In addition, we introduce CWEM, a comprehensive evaluation metric that evaluates watermarking techniques based on correctness, detectability, and naturalness. While correctness and detectability have been widely used, naturalness remains underexplored despite its importance. Unnatural patterns can reveal the presence of a watermark, making it easier for adversaries to remove. We evaluate STONE using CWEM and compare its performance with the state-of-the-art approach. The results show that STONE achieves an average improvement of 7.69% in CWEM across Python, C++, and Java. Our code is available in https://github.com/inistory/STONE-watermarking/.
[…]
linecolor=white, backgroundcolor=gray!20,
Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code
footnotetext: $\star$ Authors equally contributed.footnotetext: $\dagger$ Corresponding author.
1 Introduction
The rapid development of large language models (LLMs) has significantly improved the ability to generate human-like text and code (Nam et al., 2024; Kazemitabaar et al., 2023; Liu et al., 2023a; Sun et al., 2024; Guo et al., 2024a; Wang et al., 2024; Wei et al., 2023; Tang et al., 2023; Guo et al., 2024b). These advancements have unlocked new possibilities across various fields, including software development and automated code generation. However, this progress has introduced challenges in tracing the origin of generated code (Yang et al., 2023; Li et al., 2023). As LLM-generated code becomes increasingly indistinguishable from human-written code, researchers have explored watermarking techniques to address this issue.
LLM watermarking involves embedding specific patterns into the generated output during the generation process to help identify its origin (Kirchenbauer et al., 2023; Lee et al., 2024). These patterns should be imperceptible to human reviewers but remain detectable through algorithmic analysis, enhancing the transparency and accountability of AI-generated content (Guo et al., 2024c; Hou et al., 2024).
Among various approaches, the green and red list-based watermarking is the most widely adopted (Kirchenbauer et al., 2023; Zhao et al., 2023; Lee et al., 2024; Lu et al., 2024; Guo et al., 2024c). This method divides token candidates into two lists and biases the selection process toward tokens in the green list. Detection involves measuring the proportion of green list tokens in the generated output. While effective for natural language tasks, this approach does not easily apply to code generation, where structural and functional correctness are crucial (Lee et al., 2024; Hoang et al., 2024). Changes to syntax-critical tokens can cause compilation errors or alter program behavior.
Code watermarking identifies AI-generated code by embedding patterns into the code while maintaining functional correctness. Effective watermarking requires meeting two conditions: reliable watermark detectability and intact code correctness. However, many existing methods inadvertently modify tokens essential to program logic, such as keywords in conditional expressions or operators in arithmetic computations. These modifications can introduce syntax errors or change the behavior of the code, undermining its reliability. We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively inserts watermarks into non-syntax tokens. By excluding tokens that are critical to code execution, STONE reduces the risk of functional degradation while embedding robust, detectable patterns.
Additionally, we introduce CWEM (Code Watermarking Evaluation Metric), a comprehensive framework for evaluating watermarking techniques. CWEM assesses performance across three key dimensions: correctness, detectability, and naturalness. While correctness and detectability have been the focus of prior research, naturalness is equally important. Watermarking methods that create unnatural code patterns make it easier for adversaries to detect and remove the embedded patterns. We evaluate STONE using CWEM and compare its performance with the state-of-the-art approach. Our experiments demonstrate that STONE achieves an average improvement of 7.69% in CWEM across Python, C++, and Java.
2 Preliminary analysis
The previous SOTA approach SWEET (Lee et al., 2024) takes into account the structural and syntactic constraints by embedding watermarks only in high-entropy tokens, it ensures a minimal impact on the code quality. In our preliminary analysis, we group code tokens into several categories and compare their average entropy values to see if high-entropy tokens have a smaller impact on code quality. We consider two popular benchmarks, MBPP+ and HumanEval+, for the Python code generation and completion. We classify each Python code into one of five categories reported in Table 7 where any remaining tokens are put into the etc. category.
We calculate the entropy values of tokens across different categories and compare the average entropy values for each category. Following Lee et al. (2024), we calculate the Shannon entropy. Given a token sequence $y=(y_{0},y_{1},\ldots,y_{N})$, the entropy at each generation step $t$ is computed based on the predicted probability distribution of the next token. The entropy $H_{t}$ is defined as:
$H_{t}=-\sum_{i=1}^{|V|}P(y_{t}=v_{i}\mid y_{<t})\log P(y_{t}=v_{i}\mid y_{<t}),$ (1)
where $H_{t}$ is the entropy value at generation step $t$, $V$ is the vocabulary of the language model, $y_{<t}$ is the sequence of tokens generated before step $t$, and $P(y_{t}=v_{i}\mid y_{<t})$ is the probability assigned by the model to token $v_{i}$ as the next token, respectively. The entropy $H_{t}$ measures the uncertainty of the next token prediction. A high entropy value indicates that the model is uncertain and many tokens have comparable probabilities. Conversely, low entropy implies that one token has a dominant probability.
Figure 1: Entropy values by token types. We refer to tokens other than keywords, whitespaces, types, delimiters, and operators as etc tokens.
Figure 1 presents bar charts displaying the mean entropy values for each token category. Our analysis shows that the average entropy of keyword tokens is the highest among all categories. In Python code, keywords are used in various contexts and positions. For example, control flow keywords like if, for, and while can appear in different parts of the code with highly diverse usage patterns. In contrast, operators and delimiters follow more predictable patterns due to the syntactic structure of the code. For instance, operators such as +, -, and * are used between operands, while delimiters like ( and ) appear in specific patterns, such as in function or class definitions. Additionally, type tokens are mainly used in limited contexts, such as variable declarations or type hinting, making their occurrence and patterns relatively predictable. Whitespace tokens, following the structure of the code, are even more predictable in their placement.
Since keywords have high entropy, embedding watermarks in high-entropy tokens, as in SWEET, carries the risk of placing watermarks on keywords. Keywords are essential to the syntax and logic of the code, thus watermarking them could negatively impact the quality and functionality of the code. In comparison, the etc category, while less critical to the functionality of the code, has relatively high entropy. Therefore, embedding watermarks in etc tokens may be a more suitable approach for preserving code quality.
Figure 2: A motivation example comparing SWEET (Lee et al., 2024) and STONE (ours). (a) An example prompt for code generation, (b) Solution code without any watermark, (c) Output code watermarked by SWEET, and (d) Output code watermarked by STONE. In this example, SWEET accidentally replaces % (modulus) with ///// / (integer division) when checking if a number is even, leading to incorrect logic.
3 Approach: STONE
STONE watermarking is a robust watermarking approach designed for preserving code quality illustrated in Figure 2. This section introduces the watermark insertion and detection process of our proposed method. In Section 3.1, we detail how watermarks are inserted by selectively modifying non-syntactic tokens while keeping critical elements like keywords, operators, and delimiters unchanged to preserve functionality and structure. In Section 3.2, we explain how the watermark is detected by analyzing the green token distribution among non-syntax elements and computing a statistical $z$-score to confirm the presence of the watermark.
3.1 Inserting Watermarks
Algorithm 1 Generation Algorithm of STONE
1:Tokenized prompt $\mathbf{x}$, previously generated tokens $\mathbf{y}_{:t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
2:for $t=0,1,2,\dots$ do
3:     Compute the logit vector: $\mathbf{l}_{t}=f_{\text{LM}}(\mathbf{x},\mathbf{y}_{:t})$
4:     Compute the initial probability vector $\mathbf{p}_{t}$:
$p_{t,i}=\frac{e^{l_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{l_{t}[j]}},\quad\text{for }i\in\{0,\dots,|\mathcal{V}|-1\}$
5:     Sample a candidate token $\tilde{y}_{t}\sim\mathbf{p}_{t}$
6:     if $\tilde{y}_{t}\notin S$ then
7:         Compute a hash of token $y_{t-1}$
8:         Randomly divide $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
$\mathcal{G}_{t}\quad(\text{green list of size }\gamma|\mathcal{V}|)$
$\mathcal{R}_{t}\quad(\text{red list of size }(1-\gamma)|\mathcal{V}|)$
9:         Add $\delta$ to the logits of tokens in $\mathcal{G}_{t}$:
$l_{t}[i]\leftarrow l_{t}[i]+\delta,\quad\text{for }i\in\mathcal{G}_{t}$
10:         Recompute the adjusted probability vector $\mathbf{p}^{\prime}_{t}$:
$p^{\prime}_{t,i}=\frac{e^{l_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{l_{t}[j]}},\quad\text{for }i\in\{0,\dots,|\mathcal{V}|-1\}$
11:     else
12:         Set $\mathbf{p}^{\prime}_{t}=\mathbf{p}_{t}$
13:     end if
14:     Sample the final token $y_{t}\sim\mathbf{p}^{\prime}_{t}$
15:end for
We define the keywords, whitespaces, types, delimiters, and operators presented in Table 7 as functionally important syntactic elements and embed watermarks only in tokens that do not belong to these categories. Algorithm 1 describes how STONE embeds watermarks during code generation. Given a tokenized prompt $\mathbf{x}$ and $\mathbf{y}_{:t}$, the tokens generated up to the previous time step, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ for the token to be generated at the current time step $t$. STONE samples a candidate token $\tilde{y}_{t}$ based on this initial probability distribution. If the sampled token is not in the syntax element set, STONE divides the vocabulary $\mathcal{V}$ into a green list and a red list according to the green token ratio $\gamma$. It then increases the likelihood of sampling tokens from the green list by adding a constant $\delta$ to their logit values. Using the adjusted logits, STONE recalculates the probability distribution and samples the final token $y_{t}$ for time step $t$. Since STONE embeds watermarks while preserving functionally important tokens, it is able to maintain the quality of the code.
3.2 Detecting Watermarks
Algorithm 2 Detection Algorithm of STONE
1:Token sequence to be tested $\mathbf{X}=(X_{0},\dots,X_{N-1})$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, $z$-score threshold $z_{\text{threshold}}>0$
2:Set $N^{E}\leftarrow 0$, $N^{E}_{G}\leftarrow 0$
3:for $t=0,1,\dots,N-1$ do
4:     if $X_{t}\notin S$ then $\triangleright$ Non-syntax element detected
5:         $N^{E}\leftarrow N^{E}+1$
6:         Compute a hash of token $X_{t-1}$ to use as a seed
7:         Split vocabulary $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
$\mathcal{G}_{t}\quad(\text{green list of size }\gamma|\mathcal{V}|)$
$\mathcal{R}_{t}\quad(\text{red list of size }(1-\gamma)|\mathcal{V}|)$
8:         if $X_{t}\in\mathcal{G}_{t}$ then
9:              $N^{E}_{G}\leftarrow N^{E}_{G}+1$triangleright Token matches the green list
10:         end if
11:     end if
12:end for
13:Compute the $z$-score: $z=\frac{N^{E}_{G}-\gamma N^{E}}{\sqrt{\gamma(1-\gamma)N^{E}}}$
14:if $z>z_{\text{threshold}}$ then
15:     return True $\triangleright$ The sequence $\mathbf{X}$ is watermarked
16:else
17:     return False $\triangleright$ The sequence $\mathbf{X}$ is not watermarked
18:end if
We assess how many green tokens are present in the code, and if this value surpasses a certain threshold, we conclude that the code was generated by an LLM. When identifying potential green tokens, we focus only on non-syntactic elements (i.e., etc tokens). Algorithm 2 details the watermark detection process for STONE scenario. $N^{E}$ represents the number of etc tokens in the code, and $N^{E}_{G}$ denotes the number of green tokens among them. We replicate the division of the vocabulary into green and red lists at each time step during code generation to determine whether a specific token is a green token. If the $z$-score exceeds the predefined threshold $z_{\text{threshold}}$, the code is determined to be generated by an LLM. The purpose of embedding and detecting watermarks is to determine whether a given piece of code was generated by an LLM.
4 Code Watermarking Evaluation Metric
Existing code watermarking research lacks a unified evaluation metric that simultaneously considers functional correctness, detectability and naturalness (Lee et al., 2024; Yang et al., 2024; Li et al., 2024; Guan et al., 2024). They have typically focused one or two of these properties, leading to incomplete assessments and conflicting interpretations. By integrating functional correctness, detectability, and naturalness, we propose CWEM (Code Watermarking Evaluation Metric), a flexible metric that balances each factor using weights $\alpha$,$\beta$ and $\gamma$ satisfying $\alpha$ + $\beta$ + $\gamma$ = 1. In practice, these coefficients reflect the relative importance of functional correctness ($\alpha$ ), watermark detectability ($\beta$) and fluency or naturalness ($\gamma$). For instance, systems that cannot risk functional errors can assign a larger $\alpha$, while environments prioritizing watermark detection may emphasize $\beta$.
[…]
$+\gamma\cdot\mathrm{Naturalness}(C_{wm},C),$
correctness is measured through the pass@$k$ change before and after watermarking (Section 4.1), detectability is quantified using $z$-score and AUROC metric (Section 4.2) and naturalness is assessed using perplexity (Section 4.3).
4.1 Correctness
Ensuring correctness is crucial to prevent watermarking from compromising the functionality of the code. We measure functionality preservation using the pass@k metric. Specifically, we adopt the unbiased version of pass@$k$ (Chen et al., 2021). In this approach, $k$ samples per problem are generated, and the number of correct samples, $c$, is counted as the number of watermarked code samples in the set ${C_{wm}}$ that pass all unit tests.
This metric estimates pass@$k$ with the formula:
$\mathrm{Correctness}(C_{wm})=\mathbb{E}_{C_{wm}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right],$ (3)
where $\mathbb{E}_{C_{wm}}$ denotes the expectation over the set of watermarked code samples $C_{wm}$, $n$ is the total number of generated code samples for a given problem, $c$ is the number of correct code samples within the set, $\binom{n}{k}$ denotes the number of ways to choose $k$ code samples from the total $n$ samples, and $\binom{n-c}{k}$ represents the number of ways to choose $k$ code samples from the incorrect samples only (i.e., excluding the $c$ correct samples). The term $1-\frac{\binom{n-c}{k}}{\binom{n}{k}}$ calculates the probability that at least one correct code sample is selected when $k$ samples are drawn from the total $n$ samples. The overall correctness metric is obtained as the expected value of this probability over the set $C_{wm}$.
4.2 Detectability
Detectability measures how effectively we can identify the presence of a watermark in generated code. We assess the detectability by computing a $z$-score based on the distribution of green-list tokens in the generated text. The classification performance is then evaluated using AUROC (Area Under the Receiver Operating Characteristic Curve).
$Z$-score: Green Token Ratio
For a given code token sequence $C_{wm}=\{c_{0},c_{1},...,c_{T}\}$, the $z$-score is computed as:
$z(C_{wm})=\frac{|C_{wm}|_{G}-\gamma T}{\sqrt{T\gamma(1-\gamma)}},$ (4)
where $|C_{wm}|_{G}$ is the number of green-list tokens in the sequence $C_{wm}$, $T$ is the total number of tokens in the sequence, and $\gamma$ is the expected proportion of green tokens in a non-watermarked sequence. The denominator represents the standard deviation under the binomial distribution assumption. This $z$-score quantifies the deviation of the observed green token frequency from its expected value, enabling the detection of watermarked text.
LLM-Generated Code Detection Using Green Token Ratio
We evaluate classification performance using AUROC (Area Under the Receiver Operating Characteristic Curve) , which quantifies how well we can distinguish between human-written code $C_{H}$ and watermarked code $C_{wm}$ based on their $z$-scores.
We treat $z(C_{H})$ as scores for the negative class (human-written) and $z(C_{wm})$ as scores for the positive class (watermarked). The ROC curve is obtained by varying a decision threshold $\tau$ computing the corresponding True Positive Rate (TPR) and False Positive Rate (FPR). TPR is the fraction of watermarked sequences correctly identified as watermarked.
$\text{TPR}(\tau)=\frac{\sum_{j}\mathbf{1}(z(C_{wm}^{(j)})>\tau)}{|{C_{wm}}|},$ (5)
FPR is the fraction of human-written sequences incorrectly classified as watermarked.
$\text{FPR}(\tau)=\frac{\sum_{i}\mathbf{1}(z(C_{H}^{(i)})>\tau)}{|{C_{H}}|},$ (6)
For each threshold $\tau$, a sequence is classified as watermarked if its computed $z$-score exceeds $\tau$. We define overall detectability across all possible thresholds as follows:
[…]
This integral computes the AUROC, which quantifies the ability of the watermarking method to differentiate between human-written and watermarked code. A higher AUROC value (closer to 1) indicates stronger detectability, whereas an AUROC of 0.5 implies that the distributions of human-written and watermarked code are indistinguishable.
4.3 Naturalness
A watermark that significantly alters token distributions can be identified and removed through pre-processing. We adopt perplexity as a direct measure of token-level fluency, which aligns closely with how language models predict code tokens. Other code-style metrics exist, but we rely on perplexity because it has been widely recognized in language modeling research as a reliable indicator of naturalness. A well-designed watermark minimizes perplexity shifts, ensuring that watermarked code remains indistinguishable from non-watermarked code.
For a given set of watermarked code sequences $C_{wm}$, perplexity is computed as:
$\text{PPL}(C_{wm})=\frac{1}{|C_{wm}|}\sum_{j=1}^{|C_{wm}|}\exp\left(-\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\log P(w_{i}^{(j)}|w_{<i}^{(j)})\right),$ (8)
where $N_{j}$ is the number of tokens in the j-th sample of $C_{wm}$, and $P(w_{i}^{(j)}|w_{<i}^{(j)})$ represents the probability of token $w_{i}^{(j)}$ given its preceding context. An effectively concealed watermark minimize perplexity shifts, ensuring imperceptibility. We define the naturalness metric as:
$\mathrm{Naturalness}(C_{wm},C)=1-\frac{|\text{PPL}(C_{wm})-\text{PPL}(C)|}{\text{PPL}(C)},$ (9)
This metric quantifies the perplexity-based difference between LLM-generated code with $C_{wm}$ and without $C$ a watermark. The metric normalizes the absolute perplexity difference by dividing it by the perplexity of non-watermarked code $C$, ensuring invariance. A value of 1 indicates that $C$ and $C_{wm}$ have identical perplexities, implying minimal divergence despite watermark insertion. Conversely, a value close to 0 suggests substantial perplexity divergence, indicating a significant impact of watermark insertion on code generation. By scaling the perplexity difference between 0 and 1, this metric provides a clear and interpretable measure of similarity.
Based on these three criteria, we propose CWEM, a comprehensive evaluation metric for assessing code watermarking effectiveness. Our metric not only enables comparison across different watermarking methods but also offers flexibility through its weighted structure to adjust varying requirements in different application contexts.
5 Results and Analysis
5.1 Experiment Settings
Datasets
For our main experiments, we use the HumanEval+ and MBPP+ datasets (Liu et al., 2023b) for Python code completion and generation tasks, following prior work (Lee et al., 2024). In addition, we incorporate the HumanEvalPack (Muennighoff et al., 2024) to evaluate the performance of our approach across three programming languages: Python, C++, Java. Table 1 shows the data statistics of the datasets we used in our experiments.
Datasets Problems Test Cases Avg Tokens
MBPP+ 399 105 87
HumanEval+ 164 764 57
HumanEvalPack-C++ 164 764 100
HumanEvalPack-Java 164 764 97
Table 1: Data statistics of datasets, including the number of problems and average test cases per problem and average tokens of human-written solution codes.
Baselines
We consider the following three baselines for comparison. (1) KGW (Kirchenbauer et al., 2023) generates text by dividing the vocabulary into a green and red list at each time step of token generation, increasing the probability of generating tokens from the green list. (2) EWD (Lu et al., 2024) leverages entropy-based token weighting during detection, giving more influence to high-entropy tokens, thus enhancing detection in texts with varying entropy distributions. (3) SWEET (Lee et al., 2024) selectively embeds watermarks in high-entropy tokens, addressing detection challenges posed by the low entropy in code. We exclude CodeIP from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods.
Method MBPP+ HumanEval+
Correctness Detectability Naturalness CWEM Correctness Detectability Naturalness CWEM
KGW 0.499 0.831 0.994 0.775 0.573 0.523 0.986 0.694
EWD 0.499 0.965 0.994 0.819 0.573 0.730 0.986 0.763
SWEET 0.502 0.867 0.992 0.787 0.574 0.710 0.978 0.754
STONE 0.571 0.982 0.990 0.848 0.587 0.777 0.978 0.781
Table 2: Experimental results on MBPP+ and HumanEval+.
Method HumanEvalPack-C++ HumanEvalPack-Java
Correctness Detectability Naturalness CWEM Correctness Detectability Naturalness CWEM
KGW 0.576 0.621 0.993 0.730 0.387 0.546 0.993 0.642
EWD 0.576 0.681 0.993 0.750 0.387 0.646 0.993 0.675
SWEET 0.584 0.641 0.979 0.735 0.413 0.580 0.901 0.631
STONE 0.622 0.729 0.990 0.780 0.445 0.721 0.979 0.715
Table 3: Experimental results on HumanEvalPack C++ and Java.
Base Model
Our base model is Qwen2.5-Coder (Hui et al., 2024), which is an advanced code generation model. According to the EvalPlus leaderboard [^1] https://evalplus.github.io/leaderboard.html, Qwen2.5-Coder ranks third, following GPT-o1 models. Given its strong performance on the leaderboard, we selected Qwen2.5-Coder as our base LLM for this study.
Evaluation Metrics
We conduct experimental evaluations using our proposed evaluation metric CWEM. The evaluation framework specifically assessed three critical aspects: functional correctness to ensure proper operation (Correctness), detectability to verify watermark identification (Detectability), and naturalness to confirm the imperceptibility of the watermark (Naturalness). We measure correctness score using the average of pass@k (k=1,5).
Implementation Details
We conduct experiments on MBPP+ and HumanEval+ under the instruction-tuned Qwen2.5-Coder 7B model. For the hyperparameter optimization, we set $\delta$=1.0, $\gamma$=0.5 for MBPP+ and $\delta$=5.0, $\gamma$=0.5 for HumanEval+ and HumanEvalPack-C++ and Java, from $\delta$=[0.5, 1.0, 2.0, 3.0, 4.0], $\gamma$=[0.25, 0.5, 0.75]. A detailed analysis of the impact of varying $\delta$ and $\gamma$ across multiple values is provided in Section C. All experiments are conducted on NVIDIA A6000 GPU.
5.2 Research Questions and Findings
We formulate three key research questions and conduct a thorough analysis to address them:
  • RQ1: How well does STONE preserve code functionality?
  • RQ2: How superior is STONE to the baseline when considering functional correctness, detectability, and naturalness comprehensively?
  • RQ3: How does the time required for watermark insertion and detection in STONE?
These three research questions guide our comprehensive analysis of STONE focusing on the following aspects: 1) Functionality Preservation: Assessing whether STONE can insert watermarks without altering the original functionality. 2) Detectability and Naturalness: Evaluating whether the watermark inserted by STONE enables reliable detection of LLM-generated code while preserving a natural and readable code appearance. 3) Performance Overhead: Measuring the computational overhead involved in watermark insertion and detection. This analysis provides an understanding of the performance of STONE across correctness, detectability, naturalness, and efficiency.
RQ1
We assess whether STONE can embed watermarks without disrupting the functionality of the original code. The experimental results in Tables 2 and 3 demonstrate that STONE improves correctness by 7.57% compared with SWEET across all benchmarks. A detailed token-level analysis (Appendix A) reveals that entropy-based token selection of SWEET inadvertently targets syntax-related elements—such as delimeters (49.29), whitespaces (38.17%), keywords (9.44%), types (3.17%), and operators (2.78%)—which are critical for preserving functionality. By contrast, STONE selectively avoids altering these syntactic tokens, thereby preventing disruptions such as parsing errors or incorrect control flows.
Table 6 provides examples of syntax tokens used as watermark tokens in SWEET. These examples show that modifying such tokens can disrupt code functionality. For example, changing keywords like ‘True’ or ‘False’ affects boolean logic in conditionals, which may lead to incorrect branching. Altering delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can trigger parsing errors. Replacing ‘None’ can interfere with object initialization and function returns, causing failures. Modifications to operators such as ‘=’, ‘-’, or ‘%’ can yield incorrect outputs. STONE avoids these issues by leaving syntax elements unchanged, thereby preserving functional correctness in the generated code.
Method MBPP+ HumanEval+ HumanEvalPack-C++ HumanEvalPack-Java
Generation Detection Generation Detection Generation Detection Generation Detection
No watermark 3329 - 1261 - 1276 - 1416 -
KGW 3320 12.90 1268 4.19 1308 8.01 1506 8.95
EWD 3320 100.43 1268 34.51 1308 56.85 1506 59.84
SWEET 3300 100.94 1270 34.68 1454 58.08 1480 59.22
STONE 3266 13.27 1277 4.62 1300 8.48 1459 9.24
Table 4: Comparison generation and detection time performance in different watermark methods.
RQ2
STONE achieves the highest detection accuracy while maintaining code naturalness. KGW adjusts token selection across all tokens, resulting in small differences between watermarked and non-watermarked code that lower detection performance. Although EWD builds on KGW with entropy-based weighting, it still applies watermarking to every token, which leads to similar issues. In contrast, STONE selectively embeds watermarks only in non-syntax elements, ensuring that the watermark signal remains distinct and unambiguous. This ensures improved detection performance.
KGW and EWD report high naturalness scores by evenly adjusting token probabilities. In comparison, STONE keeps the code syntax unchanged and adjusts only non-syntax tokens. This strategy allows watermark embedding without altering the code structure or functionality, making STONE more natural than SWEET. SWEET modifies high-entropy tokens—including syntax elements such as keywords and operators—which shifts token probabilities and eases watermark detection.
SWEET achieves naturalness scores similar to STONE in MBPP+ and HumanEval+, likely because Python syntax tolerates small changes. However, modifications of syntax elements may create detectable patterns in other more complex programming languages like C++ and Java and affect readability. By preserving syntax and embedding watermarks only in non-syntax tokens, STONE minimizes changes in token probabilities while maintaining readability. Overall, STONE improves watermark classification without compromising code naturalness.
RQ3
We analyze the computational overhead of watermarking by comparing generation time and detection time. Generation time refers to the duration of producing watermarked code, while detection time measures how long it takes to verify a watermark. Table 4 summarizes the results for each baseline. All methods show similar generation times, which indicates that watermark insertion does not significantly delay code generation.
In contrast, STONE detects watermarks on average 86% faster than SWEET and EWD. This difference arises from the detection process. Both SWEET and EWD perform entropy calculations that require computing token probability distributions, increasing detection overhead. STONE avoids these computations by verifying the watermark through a predefined green list of non-syntax elements. As a result, STONE reduces detection time while maintaining high detectability.
6 Related Work
Research on code watermarking remains limited due to the unique challenges posed by structural constraints in code, particularly in maintaining functional correctness, detectability, and naturalness. Unlike natural language, code must adhere to strict syntactic and semantic rules, making the development of effective watermarking techniques more complex.
Lee et al. (2024) proposed SWEET, a method that selectively embeds watermarks in high-entropy tokens. While SWEET demonstrated improvements in detection capability and code quality, our preliminary analysis revealed that approximately 12.6% of these high-entropy tokens correspond to syntax elements such as reserved keywords. Modifying these critical tokens can disrupt program logic, resulting in degraded functional correctness.
Guan et al. (2024) introduced CodeIP, a grammar-guided watermarking approach by utilizing type prediction to maintain syntactic validity. This method ensured that the generated code remained grammatically correct. However, the reliance on type prediction requires significant computational resources, limiting its practicality for large-scale code generation tasks.
We address the limitations observed in SWEET and CodeIP by explicitly skipping watermark insertion for syntax-critical tokens. Our approach ensures that essential structural elements, like keywords and operators, remain unchanged, thereby preserving functional correctness without additional computational overhead.
ACW (Li et al., 2024) and SrcMarker (Yang et al., 2024) adopted watermarking techniques based on code transformations, applying semantic-preserving modifications and variable name substitutions. Although these methods maintained program functionality, the introduced changes often result in unnatural code patterns, making the watermark more susceptible to removal by adversaries.
7 Conclusion
We propose STONE, a syntax-aware watermarking method that preserves the original functionality of code. Unlike existing methods that apply watermarks without considering syntax tokens—often causing syntax errors or functional failures—STONE restricts watermarking to non-syntax tokens, maintaining code structure and enhancing detection reliability. Additionally, we introduce CWEM, a comprehensive evaluation metric that assesses correctness, detectability, and naturalness. While correctness and detectability are well-studied, naturalness has been largely overlooked. By incorporating naturalness into the evaluation, CWEM provides a more balanced assessment of watermarking techniques. Experimental results show that STONE consistently outperforms existing methods across three programming languages, achieving higher CWEM scores. These findings demonstrate that syntax-aware watermarking improves detection performance while preserving code functionality, making it more practical for real-world applications.
Limitations
This study introduces the STONE method, which preserves key functional tokens in code to maintain correctness. However, the current approach is not specifically optimized to enhance the naturalness of code during watermark insertion. Experimental results indicate that the naturalness score is lower than those achieved with methods such as KGW and EWD. Future work should explore refined token selection strategies to address this aspect.
In our study, CWEM does not incorporate a measure for evaluating the persistence of the watermark when modifications, refactoring, or token substitutions alter the code. Future investigations should develop quantitative measures to assess watermark robustness under these conditions.
We conduct experiments using the Qwen2.5-Coder model across three programming languages: Python, C++, and Java. The evaluation does not extend to additional large language model architectures, training environments, or parameter scales. Future studies should broaden the evaluation to encompass other model architectures, training strategies, and parameter scales in order to develop the scope and limitations of the approach.
Appendix A Syntax Elements Coverage in Watermarking Tokens of SWEET
In Table 5, we analyzed the SWEET methodology impact on token selection for watermarking using two metrics: (1) the proportion of tokens exceeding the entropy threshold among all tokens and (2) the proportion of syntax tokens among watermarking-selected tokens. Using SWEET optimal settings ($\gamma$=0.25, $\delta$=3.0, entropy threshold = 0.9), we examined threshold levels of [0.7, 0.8, 0.9, 1.0, 1.1]. The results show an inverse relationship between threshold levels and token selection. Lower thresholds increase overall token selection while incorporating more syntax tokens, whereas higher thresholds reduce syntax token inclusion. These findings validate their selective watermarking approach, as it effectively targets high-entropy regions while preserving syntactic structures critical to text integrity.
However, even at higher entropy thresholds, a non-negligible proportion of syntax tokens is still selected for watermarking. For the HumanEval+ dataset specifically, syntax tokens constituted 12.6% of watermarking tokens, with delimiters at 49.29%, whitespaces at 38.17%, keywords at 9.44%, types at 3.17%, and operators at 2.78%. This finding demonstrates that even in optimal settings, syntax tokens are inevitably selected as watermarking tokens. This distribution indicates that delimiters make up the majority of syntax tokens selected for watermarking, which can directly impact code execution. Table 6 illustrates the specific syntax tokens comprising the 12.6% identified in the HumanEval+ dataset. Delimiters, such as colons (‘:’) and commas (‘,’), are crucial for structuring code, and their modification can introduce syntax errors (e.g., omitting a colon in ‘for i in range(10):’ causes a parsing failure). While keywords and operators are selected less frequently, their modification can significantly alter program logic, as seen in cases like replacing ‘True’ with ‘False’ or ‘+’ with ‘-’. Given these risks, refining the token selection process to minimize disruptions to code execution, such as selectively excluding critical syntax tokens, could improve the practicality of syntax-aware watermarking.
Entropy Threshold MBPP+ HumanEval+ HumanEvalPack - C++ HumanEvalPack - Java
Watermarked Syntax Watermarked Syntax Watermarked Syntax Watermarked Syntax
0.7 26.03 12.82 35.28 12.84 30.10 12.99 26.60 11.98
0.8 21.40 12.59 32.67 12.24 25.93 12.35 22.47 11.06
0.9 19.64 11.39 28.98 12.60 22.65 11.79 18.52 11.68
1.0 16.59 11.39 24.84 11.83 18.96 10.56 16.13 11.15
1.1 14.94 10.30 23.22 11.81 15.71 09.82 14.39 10.91
Table 5: Syntax elements coverage in watermarking tokens of SWEET. Watermarked is the proportion of tokens that exceeds the entropy threshold among all tokens (watermarking tokens). Syntax is the proportion of watermarking tokens that belong to syntax elements. Bolded results correspond to the optimal entropy threshold setting used in SWEET.
Category Examples
Keywords def, or, import, for, True, assert, return, is, pass, None, False, in, not, if, from
[…]
Types int, set, str
Delimiters ‘.’,:, {, ‘,’, ], [, (, ) }
Operators *, %, =, -, **, &, ^, /, +, |
Table 6: Examples of syntax tokens selected as watermarked tokens in SWEET
Appendix B Comparison of Syntax Elements in Python, C++, and Java
This section provides examples of the syntactic elements in Python, C++, and Java that were utilized in our experiments. Table 7 categorize these elements into five groups: keywords, operators, delimiters, whitespaces, and types. The selection of syntactic elements reflects the distinct characteristics of each programming language.
Category Python C++ Java
Keywords True, False, None, and, as, assert, async, await, break, class, continue, def, del, elif, else, except, finally, for, from, global, if, import, in, is, lambda, nonlocal, not, or, pass, raise, return, try, while, with, yield alignas, alignof, and, and_eq, asm, auto,bitand, bitor, break, case, catch, class, compl, concept, const, consteval, constexpr, constinit, const_cast, continue, co_await, co_return, co_yield, decltype, default, delete, do, dynamic_cast, else, enum, explicit, export, extern, false, for, friend, goto, if, inline, mutable, namespace, new, noexcept, not, not_eq, nullptr, operator, or, or_eq, private, protected, public, register, reinterpret_cast, requires, return, sizeof, static, static_assert, static_cast, struct, switch, template, this, thread_local, throw, true, try, typedef, typeid, typename, union, using, virtual, volatile, while, xor, xor_eq, override abstract, assert, break, case, catch, class, const, continue, default, do, else, enum, extends, final, finally, for, goto, if, implements, import, instanceof, interface, native, new, null, package, private, protected, public, return, static, strictfp, super, switch, synchronized, this, throw, throws, transient, try, void, volatile, while, true, false
[…]
Types int, float, complex, str, bytes, bool, list, tuple, set, dict, NoneType int, float, double, bool, char, short, long, void, unsigned, signed, size_t, ptrdiff_t, wchar_t, char8_t, char16_t, char32_t byte, short, int, long, float, double, boolean, char, String, Object
Delimiters (, ), [, ], {, }, ‘,’, :, ‘.’, ;, @, ->, … (, ), [, ], {, }, ‘,’, :, ‘.’, ;, ->, ::, … (, ), [, ], {, }, ‘,’, :, ‘.’, ;, @, ->, ::, …
[…]
Table 7: Comparison of five types of syntax elements across Python, C++, and Java.
Appendix C Trade-off Between Correct and Detect
We analyze the trade-off between pass@1 and AUROC based on watermarking strength by adjusting two parameters: $\gamma$, which controls the ratio between the green list and red list, and $\delta$, which determines how often the model selects tokens from the green list. Figure 3 shows the results in a bubble chart, revealing a clear trade-off—improving one metric often reduces the other.
The bottom-right section of each subfigure highlights larger light-blue bubbles, where high delta and low gamma values produce higher AUROC but lower pass@1. A smaller gamma shrinks the green list, limiting the token options available during generation. At the same time, a higher delta increases the likelihood of selecting tokens from the green list. This combination strengthens the watermark signal and raises AUROC but generates less natural code, reducing pass@1.
Achieving the right balance between code quality and watermark detectability requires careful tuning of gamma and delta. Maintaining code quality and correctness remains crucial when embedding watermarks, so relying on both a low gamma and a high delta together creates undesirable results. The areas shaded in green highlight points where pass@1 and AUROC achieve a proper balance. The double circles outlined in black represent the performance reported in Table 2. For the MBPP+ dataset, these points correspond to $\delta$ = 1.0 and $\gamma$ = 0.5, while for the HumanEval+ dataset, they correspond to $\delta$ = 0.5 and $\gamma$ = 0.5. The trends in the green-shaded region suggest that using a lower delta value generally results in a better balance between pass@1 and AUROC performance.
Figure 3: The tradeoff between Correct and Detect. We analyze the impact of the green list ratio $\gamma$ (indicated by marker color) and the logit adjustment factor $\delta$ (indicated by marker size). Each point represents the functional correctness of code watermarked using the STONE (Y-axis) and the performance of classifying it against human-written code (X-axis). The green-shaded area highlights points where Correct and Detect are well balanced. Double circles with an outer black ring mark the performance points reported in Table 2.

### v2

@@
 Abstract
-Code watermarking identifies AI-generated code by embedding patterns into the code during generation.
-Effective watermarking requires meeting two key conditions: the watermark should be reliably detectable, and the code should retain its original functionality.
-However, existing methods often modify tokens that are critical for program logic, such as keywords in conditional expressions or operators in arithmetic computations.
-These modifications can cause syntax errors or functional failures, limiting the practical use of watermarking.
-We present STONE, a method that preserves functional integrity by selectively inserting watermarks only into non-syntax tokens.
-By excluding tokens essential for code execution, STONE minimizes the risk of functional degradation.
-In addition, we introduce CWEM, a comprehensive evaluation metric that evaluates watermarking techniques based on correctness, detectability, and naturalness.
-While correctness and detectability have been widely used, naturalness remains underexplored despite its importance.
-Unnatural patterns can reveal the presence of a watermark, making it easier for adversaries to remove.
-We evaluate STONE using CWEM and compare its performance with the state-of-the-art approach.
-The results show that STONE achieves an average improvement of 7.69% in CWEM across Python, C++, and Java.
-Our code is available in https://github.com/inistory/STONE-watermarking/.
-[…]
-linecolor=white, backgroundcolor=gray!20,
+Identifying LLM-generated code through watermarking poses a challenge in preserving functional correctness.
+Previous methods rely on the assumption that watermarking high-entropy tokens effectively maintains output quality.
+Our analysis reveals a fundamental limitation of this assumption: syntax-critical tokens such as keywords often exhibit the highest entropy, making existing approaches vulnerable to logic corruption.
+We present STONE, a syntax-aware watermarking method that embeds watermarks only in non-syntactic tokens and preserves code integrity.
+For its rigorous assessment, we also introduce STEM, a comprehensive framework that balances three critical dimensions: correctness, detectability, and imperceptibility.
+Across Python, C++, and Java, STONE preserves correctness, sustains strong detectability, and achieves balanced performance with minimal overhead.
+Our implementation is available at https://anonymous.4open.science/r/STONE-watermarking-AB4B/.
 Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code
+  Jungin Kim   Shinwoo Park   Yo-Sub Han Yonsei University, Seoul, Republic of Korea jungin.kim@yonsei.ac.kr, pshkhh@yonsei.ac.kr, emmous@yonsei.ac.kr
 footnotetext: $\star$ Authors equally contributed.footnotetext: $\dagger$ Corresponding author.
 1 Introduction
-The rapid development of large language models (LLMs) has significantly improved the ability to generate human-like text and code (Nam et al., 2024; Kazemitabaar et al., 2023; Liu et al., 2023a; Sun et al., 2024; Guo et al., 2024a; Wang et al., 2024; Wei et al., 2023; Tang et al., 2023; Guo et al., 2024b).
+The rapid development of LLMs has significantly improved the ability to generate human-like text and code (Nam et al., 2024; Kazemitabaar et al., 2023; Liu et al., 2023a; Sun et al., 2024; Guo et al., 2024a; Wang et al., 2024; Wei et al., 2023; Tang et al., 2023; Guo et al., 2024b).
 These advancements have unlocked new possibilities across various fields, including software development and automated code generation.
-However, this progress has introduced challenges in tracing the origin of generated code (Yang et al., 2023; Li et al., 2023).
+However, this progress has raised new challenges in tracing the provenance of generated code (Yang et al., 2023; Li et al., 2023).
 As LLM-generated code becomes increasingly indistinguishable from human-written code, researchers have explored watermarking techniques to address this issue.
-LLM watermarking involves embedding specific patterns into the generated output during the generation process to help identify its origin (Kirchenbauer et al., 2023; Lee et al., 2024).
-These patterns should be imperceptible to human reviewers but remain detectable through algorithmic analysis, enhancing the transparency and accountability of AI-generated content (Guo et al., 2024c; Hou et al., 2024).
+As illustrated in Figure 1, existing approaches often modify syntax-critical tokens such as keywords, and this modification introduces syntax errors or alters program behavior.
+In contrast, STONE preserves these tokens and embeds watermarks only into non-syntactic elements, thereby maintaining structural and functional integrity.
+Figure 1: Motivating example of STONE watermarking. Existing methods modify syntax tokens and cause syntax errors, whereas STONE preserves syntax and embeds watermarks without breaking code structure.
+LLM watermarking involves embedding specific patterns into the generated output during the generation process to help identify its origin (Sander et al., 2024; Golowich and Moitra, 2024; Hu and Huang, 2024; Giboulot and Furon, 2024; Li et al., 2024b; Pang et al., 2024; Zhou et al., 2024; Fu et al., 2024; Panaitescu-Liess et al., 2025).
+These patterns remain imperceptible to humans and detectable through algorithmic analysis, thereby promoting greater transparency and accountability in AI-generated content (Guo et al., 2024c; Hou et al., 2024).
 Among various approaches, the green and red list-based watermarking is the most widely adopted (Kirchenbauer et al., 2023; Zhao et al., 2023; Lee et al., 2024; Lu et al., 2024; Guo et al., 2024c).
-This method divides token candidates into two lists and biases the selection process toward tokens in the green list.
+These approaches divide token candidates into two lists (green and red) and biases the selection process toward tokens in the green list.
 Detection involves measuring the proportion of green list tokens in the generated output.
@@
 Changes to syntax-critical tokens can cause compilation errors or alter program behavior.
-Code watermarking identifies AI-generated code by embedding patterns into the code while maintaining functional correctness.
-Effective watermarking requires meeting two conditions: reliable watermark detectability and intact code correctness.
-However, many existing methods inadvertently modify tokens essential to program logic, such as keywords in conditional expressions or operators in arithmetic computations.
-These modifications can introduce syntax errors or change the behavior of the code, undermining its reliability.
-We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively inserts watermarks into non-syntax tokens.
+We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively embeds watermarks in non-syntax tokens.
 By excluding tokens that are critical to code execution, STONE reduces the risk of functional degradation while embedding robust, detectable patterns.
-Additionally, we introduce CWEM (Code Watermarking Evaluation Metric), a comprehensive framework for evaluating watermarking techniques.
-CWEM assesses performance across three key dimensions: correctness, detectability, and naturalness.
-While correctness and detectability have been the focus of prior research, naturalness is equally important.
-Watermarking methods that create unnatural code patterns make it easier for adversaries to detect and remove the embedded patterns.
-We evaluate STONE using CWEM and compare its performance with the state-of-the-art approach.
-Our experiments demonstrate that STONE achieves an average improvement of 7.69% in CWEM across Python, C++, and Java.
-2 Preliminary analysis
-The previous SOTA approach SWEET (Lee et al., 2024) takes into account the structural and syntactic constraints by embedding watermarks only in high-entropy tokens, it ensures a minimal impact on the code quality.
+Additionally, we introduce STEM (Summative Test framework for Evaluating code-waterMarking).
+Existing studies on code watermarking rely on separate evaluation criteria, using different metrics and measurement protocols across methods, hindering fair and consistent comparison of performance.
+STEM integrates correctness, detectability, and imperceptibility into a unified framework, providing a fair basis for evaluating overall performance while allowing task-specific weighting when needed.
+We employ STEM to evaluate STONE and compare it with state-of-the-art watermarking methods.
+Experiments across Python, C++, and Java show that STONE preserves functional correctness, achieves balanced performance across all evaluation criteria, and consistently outperforms prior methods in efficiency.
+Figure 2: Entropy values by token type.
+Tokens that do not fall under keywords, whitespace, types, delimiters, or operators are categorized as etc tokens.
+The etc category is the primary target for our proposed STONE watermarking method.
+Token entropy is measured using Qwen2.5-Coder-7B.
+2 Motivation and Preliminary Analysis
+The recent code watermarking method SWEET (Lee et al., 2024) accounts for structural and syntactic constraints by embedding watermarks only in high-entropy tokens, thereby minimizing their impact on code quality.
 In our preliminary analysis, we group code tokens into several categories and compare their average entropy values to see if high-entropy tokens have a smaller impact on code quality.
 We consider two popular benchmarks, MBPP+ and HumanEval+, for the Python code generation and completion.
-We classify each Python code into one of five categories reported in Table 7 where any remaining tokens are put into the etc.
+We classify each Python code into one of five categories (keywords, whitespace, types, delimiters, and operators) where any remaining tokens are put into the etc.
 category.
+A detailed explanation is provided in Appendix A.
 We calculate the entropy values of tokens across different categories and compare the average entropy values for each category.
@@
 $H_{t}=-\sum_{i=1}^{|V|}P(y_{t}=v_{i}\mid y_{<t})\log P(y_{t}=v_{i}\mid y_{<t}),$ (1)
-where $H_{t}$ is the entropy value at generation step $t$, $V$ is the vocabulary of the language model, $y_{<t}$ is the sequence of tokens generated before step $t$, and $P(y_{t}=v_{i}\mid y_{<t})$ is the probability assigned by the model to token $v_{i}$ as the next token, respectively.
+where $H_{t}$ is the entropy value at generation step $t$, $V$ is the vocabulary of the language model, $y_{<t}$ is the sequence of tokens generated before step $t$, and $P(y_{t}=v_{i}\mid y_{<t})$ is the probability assigned by the model to token $v_{i}$ as the next token.
 The entropy $H_{t}$ measures the uncertainty of the next token prediction.
@@
 Conversely, low entropy implies that one token has a dominant probability.
-Figure 1: Entropy values by token types. We refer to tokens other than keywords, whitespaces, types, delimiters, and operators as etc tokens.
 Figure 2 presents bar charts displaying the mean entropy values for each token category.
@@
 In comparison, the etc category, while less critical to the functionality of the code, has relatively high entropy.
-Therefore, embedding watermarks in etc tokens may be a more suitable approach for preserving code quality.
-Figure 2: A motivation example comparing SWEET (Lee et al., 2024) and STONE (ours).
-(a) An example prompt for code generation, (b) Solution code without any watermark, (c) Output code watermarked by SWEET, and (d) Output code watermarked by STONE.
-In this example, SWEET accidentally replaces % (modulus) with ///// / (integer division) when checking if a number is even, leading to incorrect logic.
-3 Approach: STONE
-STONE watermarking is a robust watermarking approach designed for preserving code quality illustrated in Figure 2.
-This section introduces the watermark insertion and detection process of our proposed method.
-In Section 3.1, we detail how watermarks are inserted by selectively modifying non-syntactic tokens while keeping critical elements like keywords, operators, and delimiters unchanged to preserve functionality and structure.
-In Section 3.2, we explain how the watermark is detected by analyzing the green token distribution among non-syntax elements and computing a statistical $z$-score to confirm the presence of the watermark.
-3.1 Inserting Watermarks
-Algorithm 1 Generation Algorithm of STONE
-1:Tokenized prompt $\mathbf{x}$, previously generated tokens $\mathbf{y}_{:t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
+Therefore, embedding watermarks in etc tokens (i.e., non-syntactic tokens) may be a more suitable approach for preserving code quality.
+3 Methodology
+3.1 Syntax-Aware Watermarking: STONE
+STONE is a code watermarking method specifically designed to preserve the quality of generated code.
+This section describes how STONE performs watermark insertion and detection.
+We describe the insertion process, where watermarks are embedded only in non-syntactic tokens while keeping critical elements such as keywords, operators, and delimiters unchanged to preserve functionality and structure.
+Detection is performed by analyzing the distribution of green tokens among non-syntax elements.
+Inserting Watermarks
+We define the keywords, whitespace, types, delimiters, and operators as functionally important syntactic elements and embed watermarks only in tokens that do not belong to these categories.
+Algorithm 1 describes how STONE embeds watermarks during code generation.
+Given a prompt $\mathbf{x}$ and $\mathbf{y}_{:t}$, the tokens generated up to the previous time step, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ for the token to be generated at the current time step $t$.
+STONE samples a candidate token $\tilde{y}_{t}$ based on this initial probability distribution.
+If the sampled token is not in the syntax element set, STONE divides the vocabulary $\mathcal{V}$ into a green/red list according to the green token ratio $\gamma$.
+It then increases the likelihood of sampling tokens from the green list by adding a constant $\delta$ to their logit values.
+Using the adjusted logits, STONE updates the probability distribution and samples the final token $y_{t}$ for time step $t$.
+Algorithm 1 Insertion Algorithm of STONE
+1:Prompt $\mathbf{x}$, previously generated tokens $\mathbf{y}_{:t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
 2:for $t=0,1,2,\dots$ do
@@
 6:  if $\tilde{y}_{t}\notin S$ then
-7:         Compute a hash of token $y_{t-1}$
-8:         Randomly divide $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
+7:   Compute a hash of token $y_{t-1}$ to use as a seed
+8:   Split $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
 $\mathcal{G}_{t}\quad(\text{green list of size }\gamma|\mathcal{V}|)$
@@
 15:end for
-We define the keywords, whitespaces, types, delimiters, and operators presented in Table 7 as functionally important syntactic elements and embed watermarks only in tokens that do not belong to these categories.
-Algorithm 1 describes how STONE embeds watermarks during code generation.
-Given a tokenized prompt $\mathbf{x}$ and $\mathbf{y}_{:t}$, the tokens generated up to the previous time step, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ for the token to be generated at the current time step $t$.
-STONE samples a candidate token $\tilde{y}_{t}$ based on this initial probability distribution.
-If the sampled token is not in the syntax element set, STONE divides the vocabulary $\mathcal{V}$ into a green list and a red list according to the green token ratio $\gamma$.
-It then increases the likelihood of sampling tokens from the green list by adding a constant $\delta$ to their logit values.
-Using the adjusted logits, STONE recalculates the probability distribution and samples the final token $y_{t}$ for time step $t$.
-Since STONE embeds watermarks while preserving functionally important tokens, it is able to maintain the quality of the code.
-3.2 Detecting Watermarks
+Detecting Watermarks
+We assess how many green tokens are present in the code, and if this value surpasses a certain threshold, we conclude that the code was generated by an LLM.
+When identifying potential green tokens, we focus only on non-syntactic elements.
+Algorithm 2 details the watermark detection process.
+$N^{E}$ represents the number of etc tokens (non-syntax tokens) in the code, and $N^{E}_{G}$ denotes the number of green tokens among them.
+We replicate the division of the vocabulary into green and red lists at each time step during code generation to determine whether a specific token is a green token.
+If the $z$-score exceeds the predefined threshold $z_{\text{threshold}}$, the code is determined to be generated by an LLM.
 Algorithm 2 Detection Algorithm of STONE
@@
 6:   Compute a hash of token $X_{t-1}$ to use as a seed
-7:         Split vocabulary $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
+7:   Split $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
 $\mathcal{G}_{t}\quad(\text{green list of size }\gamma|\mathcal{V}|)$
@@
 8:   if $X_{t}\in\mathcal{G}_{t}$ then
-9:              $N^{E}_{G}\leftarrow N^{E}_{G}+1$triangleright Token matches the green list
+9:     $N^{E}_{G}\leftarrow N^{E}_{G}+1$ $\triangleright$ Token matches the green list
 10:   end if
@@
 18:end if
-We assess how many green tokens are present in the code, and if this value surpasses a certain threshold, we conclude that the code was generated by an LLM.
-When identifying potential green tokens, we focus only on non-syntactic elements (i.e., etc tokens).
-Algorithm 2 details the watermark detection process for STONE scenario.
-$N^{E}$ represents the number of etc tokens in the code, and $N^{E}_{G}$ denotes the number of green tokens among them.
-We replicate the division of the vocabulary into green and red lists at each time step during code generation to determine whether a specific token is a green token.
-If the $z$-score exceeds the predefined threshold $z_{\text{threshold}}$, the code is determined to be generated by an LLM.
-The purpose of embedding and detecting watermarks is to determine whether a given piece of code was generated by an LLM.
-4 Code Watermarking Evaluation Metric
-Existing code watermarking research lacks a unified evaluation metric that simultaneously considers functional correctness, detectability and naturalness (Lee et al., 2024; Yang et al., 2024; Li et al., 2024; Guan et al., 2024).
-They have typically focused one or two of these properties, leading to incomplete assessments and conflicting interpretations.
-By integrating functional correctness, detectability, and naturalness, we propose CWEM (Code Watermarking Evaluation Metric), a flexible metric that balances each factor using weights $\alpha$,$\beta$ and $\gamma$ satisfying $\alpha$ + $\beta$ + $\gamma$ = 1.
-In practice, these coefficients reflect the relative importance of functional correctness ($\alpha$ ), watermark detectability ($\beta$) and fluency or naturalness ($\gamma$).
-For instance, systems that cannot risk functional errors can assign a larger $\alpha$, while environments prioritizing watermark detection may emphasize $\beta$.
-[…]
-$+\gamma\cdot\mathrm{Naturalness}(C_{wm},C),$
-correctness is measured through the pass@$k$ change before and after watermarking (Section 4.1), detectability is quantified using $z$-score and AUROC metric (Section 4.2) and naturalness is assessed using perplexity (Section 4.3).
-4.1 Correctness
-Ensuring correctness is crucial to prevent watermarking from compromising the functionality of the code.
-We measure functionality preservation using the pass@k metric.
-Specifically, we adopt the unbiased version of pass@$k$ (Chen et al., 2021).
-In this approach, $k$ samples per problem are generated, and the number of correct samples, $c$, is counted as the number of watermarked code samples in the set ${C_{wm}}$ that pass all unit tests.
-This metric estimates pass@$k$ with the formula:
+3.2 Unified Evaluation Framework: STEM
+Prior studies on code watermarking have typically evaluated functional correctness, detectability, and imperceptibility as separate criteria (Lee et al., 2024; Yang et al., 2024; Li et al., 2024a; Guan et al., 2024).
+However, evaluating each criterion in isolation makes it difficult to capture the trade-offs between them and to identify methods with balanced performance.
+For example, a watermarking technique may enhance detectability at the expense of correctness, resulting in incomplete or conflicting assessments.
+This work addresses this limitation by introducing STEM, a balanced evaluation framework that integrates correctness, detectability, and imperceptibility into a unified score.
+STEM combines the three criteria using weights $\alpha$, $\beta$, and $\zeta$ satisfying $\alpha+\beta+\zeta=1$, enabling direct comparison of methods in terms of overall balance rather than isolated metrics.
+$\text{{STEM}}=$ $\;\alpha\cdot\mathrm{Correctness}(C_{wm})$ (2)
+$+\beta\cdot\mathrm{Detectability}(C_{wm},C_{H})$
+$+\zeta\cdot\mathrm{Imperceptibility}(C_{wm},C).$
+Correctness
+Maintaining correctness is crucial for preventing watermarking from compromising code functionality.
+Let $C_{wm}$ denote the set of watermarked code samples.
+We measure functionality preservation using the unbiased estimator of pass@k Chen et al.
+(2021):
 $\mathrm{Correctness}(C_{wm})=\mathbb{E}_{C_{wm}}\!\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$], (3)
-where $\mathbb{E}_{C_{wm}}$ denotes the expectation over the set of watermarked code samples $C_{wm}$, $n$ is the total number of generated code samples for a given problem, $c$ is the number of correct code samples within the set, $\binom{n}{k}$ denotes the number of ways to choose $k$ code samples from the total $n$ samples, and $\binom{n-c}{k}$ represents the number of ways to choose $k$ code samples from the incorrect samples only (i.e., excluding the $c$ correct samples).
-The term $1-\frac{\binom{n-c}{k}}{\binom{n}{k}}$ calculates the probability that at least one correct code sample is selected when $k$ samples are drawn from the total $n$ samples.
-The overall correctness metric is obtained as the expected value of this probability over the set $C_{wm}$.
-4.2 Detectability
-Detectability measures how effectively we can identify the presence of a watermark in generated code.
-We assess the detectability by computing a $z$-score based on the distribution of green-list tokens in the generated text.
-The classification performance is then evaluated using AUROC (Area Under the Receiver Operating Characteristic Curve).
-$Z$-score: Green Token Ratio
-For a given code token sequence $C_{wm}=\{c_{0},c_{1},...,c_{T}\}$, the $z$-score is computed as:
-$z(C_{wm})=\frac{|C_{wm}|_{G}-\gamma T}{\sqrt{T\gamma(1-\gamma)}},$ (4)
-where $|C_{wm}|_{G}$ is the number of green-list tokens in the sequence $C_{wm}$, $T$ is the total number of tokens in the sequence, and $\gamma$ is the expected proportion of green tokens in a non-watermarked sequence.
-The denominator represents the standard deviation under the binomial distribution assumption.
-This $z$-score quantifies the deviation of the observed green token frequency from its expected value, enabling the detection of watermarked text.
-LLM-Generated Code Detection Using Green Token Ratio
-We evaluate classification performance using AUROC (Area Under the Receiver Operating Characteristic Curve) , which quantifies how well we can distinguish between human-written code $C_{H}$ and watermarked code $C_{wm}$ based on their $z$-scores.
-We treat $z(C_{H})$ as scores for the negative class (human-written) and $z(C_{wm})$ as scores for the positive class (watermarked).
-The ROC curve is obtained by varying a decision threshold $\tau$ computing the corresponding True Positive Rate (TPR) and False Positive Rate (FPR).
-TPR is the fraction of watermarked sequences correctly identified as watermarked.
-$\text{TPR}(\tau)=\frac{\sum_{j}\mathbf{1}(z(C_{wm}^{(j)})>\tau)}{|{C_{wm}}|},$ (5)
-FPR is the fraction of human-written sequences incorrectly classified as watermarked.
-$\text{FPR}(\tau)=\frac{\sum_{i}\mathbf{1}(z(C_{H}^{(i)})>\tau)}{|{C_{H}}|},$ (6)
-For each threshold $\tau$, a sequence is classified as watermarked if its computed $z$-score exceeds $\tau$. We define overall detectability across all possible thresholds as follows:
-[…]
-This integral computes the AUROC, which quantifies the ability of the watermarking method to differentiate between human-written and watermarked code.
-A higher AUROC value (closer to 1) indicates stronger detectability, whereas an AUROC of 0.5 implies that the distributions of human-written and watermarked code are indistinguishable.
-4.3 Naturalness
-A watermark that significantly alters token distributions can be identified and removed through pre-processing.
-We adopt perplexity as a direct measure of token-level fluency, which aligns closely with how language models predict code tokens.
-Other code-style metrics exist, but we rely on perplexity because it has been widely recognized in language modeling research as a reliable indicator of naturalness.
-A well-designed watermark minimizes perplexity shifts, ensuring that watermarked code remains indistinguishable from non-watermarked code.
-For a given set of watermarked code sequences $C_{wm}$, perplexity is computed as:
-$\text{PPL}(C_{wm})=\frac{1}{|C_{wm}|}\sum_{j=1}^{|C_{wm}|}\exp\left(-\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\log P(w_{i}^{(j)}|w_{<i}^{(j)})\right),$ (8)
-where $N_{j}$ is the number of tokens in the j-th sample of $C_{wm}$, and $P(w_{i}^{(j)}|w_{<i}^{(j)})$ represents the probability of token $w_{i}^{(j)}$ given its preceding context.
-An effectively concealed watermark minimize perplexity shifts, ensuring imperceptibility.
-We define the naturalness metric as:
-$\mathrm{Naturalness}(C_{wm},C)=1-\frac{|\text{PPL}(C_{wm})-\text{PPL}(C)|}{\text{PPL}(C)},$ (9)
-This metric quantifies the perplexity-based difference between LLM-generated code with $C_{wm}$ and without $C$ a watermark.
-The metric normalizes the absolute perplexity difference by dividing it by the perplexity of non-watermarked code $C$, ensuring invariance.
-A value of 1 indicates that $C$ and $C_{wm}$ have identical perplexities, implying minimal divergence despite watermark insertion.
-Conversely, a value close to 0 suggests substantial perplexity divergence, indicating a significant impact of watermark insertion on code generation.
-By scaling the perplexity difference between 0 and 1, this metric provides a clear and interpretable measure of similarity.
-Based on these three criteria, we propose CWEM, a comprehensive evaluation metric for assessing code watermarking effectiveness.
-Our metric not only enables comparison across different watermarking methods but also offers flexibility through its weighted structure to adjust varying requirements in different application contexts.
-5 Results and Analysis
-5.1 Experiment Settings
-Datasets
-For our main experiments, we use the HumanEval+ and MBPP+ datasets (Liu et al., 2023b) for Python code completion and generation tasks, following prior work (Lee et al., 2024).
-In addition, we incorporate the HumanEvalPack (Muennighoff et al., 2024) to evaluate the performance of our approach across three programming languages: Python, C++, Java.
-Table 1 shows the data statistics of the datasets we used in our experiments.
-Datasets Problems Test Cases Avg Tokens
-MBPP+ 399 105 87
-HumanEval+ 164 764 57
-HumanEvalPack-C++ 164 764 100
-HumanEvalPack-Java 164 764 97
-Table 1: Data statistics of datasets, including the number of problems and average test cases per problem and average tokens of human-written solution codes.
-Baselines
-We consider the following three baselines for comparison.
+where $n$ and $c$ denote the total and correct generations per task, respectively. A higher score indicates that watermark insertion does not hinder program execution.
+Detectability
+Detectability measures how effectively a watermark can be identified in generated code.
+We assess detectability using a $z$-score that quantifies the proportion of green-list tokens and compute AUROC between human-written and watermarked corpora.
+For a generated sequence $X_{wm}=(X_{0},X_{1},\dots,X_{N-1})$, the $z$-score of green-token ratio is defined as
+$z(X_{wm})=\frac{|X_{wm}|_{G}-\gamma|X_{wm}|}{\sqrt{\gamma(1-\gamma)|X_{wm}|}},$ (4)
+where $|X_{wm}|_{G}$ is the number of green tokens in $X_{wm}$ and $\gamma$ is the expected green-token ratio in non-watermarked code.
+Higher $z$ values indicate stronger evidence that the code is watermarked.
+Given a corpus of human-written code $C_{H}=\{X^{(i)}_{H}\}$ and watermarked code $C_{wm}=\{X^{(j)}_{wm}\}$, we compute the true- and false-positive rates at threshold $\tau$:
+$\text{TPR}(\tau)$ $=\frac{\sum_{j=1}^{J}\mathbf{1}\!\bigl[z(X_{wm}^{(j)})>\tau\bigr]}{J}.$ (5)
+$\text{FPR}(\tau)$ $=\frac{\sum_{i=1}^{I}\mathbf{1}\!\bigl[z(X_{H}^{(i)})>\tau\bigr]}{I}.$ (6)
+The area under the ROC curve (AUROC) is then
+$\mathrm{Detectability}(C_{wm},C_{H})=\int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR}),$ (7)
+where higher AUROC values indicate stronger detectability between watermarked and human-written code.
+Imperceptibility
+Watermarking should preserve token probability distributions consistent with normal model outputs (Huang and Wan, 2025).
+In code watermarking, Sun et al.
+(2023) assessed imperceptibility through post-hoc detection combining human and automated evaluations, which does not quantify distributional distortion.
+Following perplexity-based approaches in text watermarking (Kirchenbauer et al., 2023) and studies including code generation (Magnusson et al., 2024), we define imperceptibility as the deviation in token probability distributions introduced by watermarking, measured through perplexity shifts computed with code LLMs.
+Let $C_{wm}$ and $C$ denote the corpora of watermarked and non-watermarked code, respectively. Perplexity is computed as
+$\operatorname{PPL}(C_{wm})=\begin{aligned} &\frac{1}{\lvert C_{wm}\rvert}\sum_{j=1}^{\lvert C_{wm}\rvert}\exp\!\bigl(-A_{j}\bigr),\\ &A_{j}=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\log P\!\left(y_{i}^{(j)}\mid y_{<i}^{(j)}\right).\end{aligned}$ (8)
+Higher values indicate that watermark insertion produces little distributional deviation from normal code generation.
+4 Experimental Setup
+Datasets.
+For our experiments, we evaluate our approach across three programming languages—Python, C++, and Java—using four benchmark datasets: HumanEval+ and MBPP+ (Liu et al., 2023b) for Python, and HumanEvalPack (Muennighoff et al., 2024) for C++ and Java.
+Table 1 presents token-length statistics of the evaluation datasets, computed with the Qwen2.5-Coder-7B tokenizer.
+The datasets span from short snippets to long sequences, enabling assessment of our method across varying complexity.
+Dataset Problems Test Cases Max Min Mean Std.
+MBPP+ 399 105 341 11 40.25 34.91
+HumanEval+ 164 764 558 43 188.28 87.32
+HEP-C++ 164 764 697 42 223.10 107.57
+HEP-Java 164 764 660 56 237.36 105.09
+Table 1: Token length statistics of solution code across evaluation datasets, along with the total number of problems and the average number of test cases per problem.
+Baselines.
+We consider the following three watermarking approaches for comparison.
 (1) KGW (Kirchenbauer et al., 2023) generates text by dividing the vocabulary into a green and red list at each time step of token generation, increasing the probability of generating tokens from the green list.
@@
 (3) SWEET (Lee et al., 2024) selectively embeds watermarks in high-entropy tokens, addressing detection challenges posed by the low entropy in code.
-We exclude CodeIP from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods.
+We exclude CodeIP (Guan et al., 2024) from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods.
+As part of our comparative analysis, we reproduce CodeIP and report its experimental results in Appendix B.
+Base Model.
+We use Qwen2.5-Coder-7B (Hui et al., 2024) for our experiments.
+We also report additional results with Llama-3.1-8B in Appendix C, which further demonstrate the generalizability of our approach across different model architectures.
+Evaluation Metric.
+We evaluate watermarking methods across three dimensions: Correctness, Detectability, and Imperceptibility.
+Correctness uses pass@k (k=1,5), detectability is measured by AUROC with a $z$-score test, and imperceptibility by perplexity computed with StarCoder2-7B.
+We use the STEM framework to integrate these three axes into a single composite score.
+For main results (Tables 2 and 3), we adopt the equal-weight configuration $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$ as a neutral.
+We further analyze weight settings with alternative configurations in Appendix D.
+Implementation Details.
+During code generation, we apply top-$k$ sampling with $k=50$, restricting the candidate token pool to the top 50 tokens with the highest logits and normalizing their probabilities before sampling.
+The temperature, which controls the sharpness of the token distribution, is fixed at 1.0.
+We set $\gamma=0.5$ and $\delta=1.0$ for the MBPP+ dataset, while for the HumanEval+ and HumanEvalPack (C++, Java) datasets, we set $\gamma=0.5$ and $\delta=0.5$.
+We conduct all experiments on an NVIDIA A6000 GPU.
 Method MBPP+ HumanEval+
-Correctness Detectability Naturalness CWEM Correctness Detectability Naturalness CWEM
+Correctness Detectability Imperceptibility STEM Correctness Detectability Imperceptibility STEM
 KGW 0.499 0.831 0.994 0.775 0.573 0.523 0.986 0.694
@@
 STONE 0.571 0.982 0.990 0.848 0.587 0.777 0.978 0.781
-Table 2: Experimental results on MBPP+ and HumanEval+.
-Method HumanEvalPack-C++ HumanEvalPack-Java
-Correctness Detectability Naturalness CWEM Correctness Detectability Naturalness CWEM
+Table 2: Experimental results on MBPP+ and HumanEval+. Metrics include correctness, detectability, imperceptibility, and the integrated STEM score under the equal-weight configuration.
+Method HEP-C++ HEP-Java
+Correctness Detectability Imperceptibility STEM Correctness Detectability Imperceptibility STEM
 KGW 0.576 0.621 0.993 0.730 0.387 0.546 0.993 0.642
@@
 STONE 0.622 0.729 0.990 0.780 0.445 0.721 0.979 0.715
-Table 3: Experimental results on HumanEvalPack C++ and Java.
-Base Model
-Our base model is Qwen2.5-Coder (Hui et al., 2024), which is an advanced code generation model.
-According to the EvalPlus leaderboard [^1] https://evalplus.github.io/leaderboard.html, Qwen2.5-Coder ranks third, following GPT-o1 models.
-Given its strong performance on the leaderboard, we selected Qwen2.5-Coder as our base LLM for this study.
-Evaluation Metrics
-We conduct experimental evaluations using our proposed evaluation metric CWEM.
-The evaluation framework specifically assessed three critical aspects: functional correctness to ensure proper operation (Correctness), detectability to verify watermark identification (Detectability), and naturalness to confirm the imperceptibility of the watermark (Naturalness).
-We measure correctness score using the average of pass@k (k=1,5).
-Implementation Details
-We conduct experiments on MBPP+ and HumanEval+ under the instruction-tuned Qwen2.5-Coder 7B model.
-For the hyperparameter optimization, we set $\delta$=1.0, $\gamma$=0.5 for MBPP+ and $\delta$=5.0, $\gamma$=0.5 for HumanEval+ and HumanEvalPack-C++ and Java, from $\delta$=[0.5, 1.0, 2.0, 3.0, 4.0], $\gamma$=[0.25, 0.5, 0.75].
-A detailed analysis of the impact of varying $\delta$ and $\gamma$ across multiple values is provided in Section C.
-All experiments are conducted on NVIDIA A6000 GPU.
-5.2 Research Questions and Findings
-We formulate three key research questions and conduct a thorough analysis to address them:
-  • RQ1: How well does STONE preserve code functionality?
-  • RQ2: How superior is STONE to the baseline when considering functional correctness, detectability, and naturalness comprehensively?
-  • RQ3: How does the time required for watermark insertion and detection in STONE?
-These three research questions guide our comprehensive analysis of STONE focusing on the following aspects: 1) Functionality Preservation: Assessing whether STONE can insert watermarks without altering the original functionality.
-2) Detectability and Naturalness: Evaluating whether the watermark inserted by STONE enables reliable detection of LLM-generated code while preserving a natural and readable code appearance.
-3) Performance Overhead: Measuring the computational overhead involved in watermark insertion and detection.
-This analysis provides an understanding of the performance of STONE across correctness, detectability, naturalness, and efficiency.
+Table 3: Experimental results on HumanEvalPack-C++ and HumanEvalPack-Java.
+Metrics include correctness, detectability, imperceptibility, and the integrated STEM score under the equal-weight configuration.
+5 Experiments and Results
+We formulate three key research questions and conduct a thorough analysis to addressing them:
+  • RQ1: Can STONE preserve functional correctness?
+  • RQ2: Does STONE achieve balanced performance across correctness, detectability, and imperceptibility as measured by STEM?
+  • RQ3: How efficient is STONE in watermark insertion and detection?
+These research questions frame our analysis of STONE along three key aspects: 1) Functionality Preservation: verifying that watermark insertion does not alter program behavior; 2) Balanced Performance under STEM: evaluating whether STONE achieves a well-balanced trade-off among correctness, detectability, and imperceptibility by the STEM metric; and 3) Performance Overhead: measuring the computational cost of watermark insertion and detection.
 RQ1
-We assess whether STONE can embed watermarks without disrupting the functionality of the original code.
-The experimental results in Tables 2 and 3 demonstrate that STONE improves correctness by 7.57% compared with SWEET across all benchmarks.
-A detailed token-level analysis (Appendix A) reveals that entropy-based token selection of SWEET inadvertently targets syntax-related elements—such as delimeters (49.29), whitespaces (38.17%), keywords (9.44%), types (3.17%), and operators (2.78%)—which are critical for preserving functionality.
-By contrast, STONE selectively avoids altering these syntactic tokens, thereby preventing disruptions such as parsing errors or incorrect control flows.
-Table 6 provides examples of syntax tokens used as watermark tokens in SWEET.
-These examples show that modifying such tokens can disrupt code functionality.
-For example, changing keywords like ‘True’ or ‘False’ affects boolean logic in conditionals, which may lead to incorrect branching.
-Altering delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can trigger parsing errors.
-Replacing ‘None’ can interfere with object initialization and function returns, causing failures.
-Modifications to operators such as ‘=’, ‘-’, or ‘%’ can yield incorrect outputs.
-STONE avoids these issues by leaving syntax elements unchanged, thereby preserving functional correctness in the generated code.
-Method MBPP+ HumanEval+ HumanEvalPack-C++ HumanEvalPack-Java
-Generation Detection Generation Detection Generation Detection Generation Detection
-No watermark 3329 - 1261 - 1276 - 1416 -
+We examine whether STONE embeds watermarks without affecting the functional correctness of generated code.
+Tables 2 and 3 show that STONE improves correctness by 7.57% over SWEET across all benchmarks.
+The token-level analysis in Appendix E indicates that the entropy-based token selection method used by SWEET often chooses syntax-related elements—such as delimiters, whitespace, keywords, types, and operators—that are essential for correct execution.
+For example, modifying delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can cause parsing errors, and changing operators or keywords (e.g., replacing ‘+’ with ‘-’ or ‘True’ with ‘False’) can alter program logic or control flow.
+Even under the optimal configuration, syntax tokens account for 12.6% of the tokens selected for watermarking in the HumanEval+ dataset, with delimiters and whitespace forming the majority.
+This inclusion of syntax elements explains the observed reduction in correctness for SWEET.
+In contrast, STONE excludes syntax tokens and embeds watermarks only into non-syntactic elements, which prevents structural and logical errors during execution and allows the generated code to preserve its functional behavior.
+RQ2
+We examine whether STONE achieves balanced performance across correctness, detectability, and imperceptibility.
+Baselines such as KGW and EWD watermark every token, and this process dilutes the watermark signal and lowers AUROC scores, while SWEET modifies high-entropy tokens including syntax elements, improving detectability but harming fluency.
+In contrast, STONE embeds watermarks only in non-syntax tokens, and this design preserves syntactic integrity while keeping the signal distinct.
+This selective design enables STONE to achieve higher detection accuracy without compromising the imperceptibility of code.
+Imperceptibility results reinforce this finding.
+KGW and EWD achieve reasonable fluency by uniformly adjusting token probabilities, but their adjustments can still introduce subtle artifacts.
+SWEET often produces more noticeable distributional shifts because it alters syntax tokens.
+STONE, on the other hand, maintains low perplexity differences by leaving syntax untouched and altering only semantically flexible positions.
+As a result, STONE sustains imperceptibility while preserving correctness and robustness in detection.
+This balanced performance is captured by the STEM score under the equal-weight configuration, which provides a neutral evaluation across correctness, detectability, and imperceptibility.
+Unlike other watermarking methods that exhibit clear trade-offs among these criteria, STONE achieves consistently superior overall performance.
+A comprehensive grid analysis across 66 weight settings (Appendix D) demonstrates that this advantage remains stable and does not depend on any specific weighting setting.
+Method MBPP+ HumanEval+ HEP-C++ HEP-Java
+Insertion Detection Insertion Detection Insertion Detection Insertion Detection
 KGW 3320 12.90 1268 4.19 1308 8.01 1506 8.95
@@
 STONE 3266 13.27 1277 4.62 1300 8.48 1459 9.24
-Table 4: Comparison generation and detection time performance in different watermark methods.
-RQ2
-STONE achieves the highest detection accuracy while maintaining code naturalness.
-KGW adjusts token selection across all tokens, resulting in small differences between watermarked and non-watermarked code that lower detection performance.
-Although EWD builds on KGW with entropy-based weighting, it still applies watermarking to every token, which leads to similar issues.
-In contrast, STONE selectively embeds watermarks only in non-syntax elements, ensuring that the watermark signal remains distinct and unambiguous.
-This ensures improved detection performance.
-KGW and EWD report high naturalness scores by evenly adjusting token probabilities.
-In comparison, STONE keeps the code syntax unchanged and adjusts only non-syntax tokens.
-This strategy allows watermark embedding without altering the code structure or functionality, making STONE more natural than SWEET.
-SWEET modifies high-entropy tokens—including syntax elements such as keywords and operators—which shifts token probabilities and eases watermark detection.
-SWEET achieves naturalness scores similar to STONE in MBPP+ and HumanEval+, likely because Python syntax tolerates small changes.
-However, modifications of syntax elements may create detectable patterns in other more complex programming languages like C++ and Java and affect readability.
-By preserving syntax and embedding watermarks only in non-syntax tokens, STONE minimizes changes in token probabilities while maintaining readability.
-Overall, STONE improves watermark classification without compromising code naturalness.
+Table 4: Comparison of total insertion and detection time (measured in seconds) across different watermarking methods.
 RQ3
-We analyze the computational overhead of watermarking by comparing generation time and detection time.
-Generation time refers to the duration of producing watermarked code, while detection time measures how long it takes to verify a watermark.
-Table 4 summarizes the results for each baseline.
-All methods show similar generation times, which indicates that watermark insertion does not significantly delay code generation.
-In contrast, STONE detects watermarks on average 86% faster than SWEET and EWD.
-This difference arises from the detection process.
-Both SWEET and EWD perform entropy calculations that require computing token probability distributions, increasing detection overhead.
-STONE avoids these computations by verifying the watermark through a predefined green list of non-syntax elements.
-As a result, STONE reduces detection time while maintaining high detectability.
+We analyze the computational overhead of watermarking by comparing insertion and detection time.
+Insertion time refers to the duration of producing watermarked code, while detection time measures how long it takes to verify a watermark.
+Table 4 presents a summary of the results for each method.
+While all approaches exhibit comparable insertion times, STONE achieves detection speeds that are, on average, 86% faster than those of SWEET and EWD.
+This improvement stems from differences in the detection process: both SWEET and EWD rely on entropy-based method, which involves computing token probability distributions and leads to greater computational overhead.
+In contrast, STONE eliminates this burden by verifying watermarks using a predefined green list of non-syntax tokens.
+This design reduces detection time without sacrificing detectability.
+6 Analysis
+6.1 Trade-off Between Code Quality and Detection Performance
+We analyze the trade-off between code quality and watermark detectability as influenced by two watermarking parameters.
+The parameter $\gamma$ controls the proportion of the vocabulary designated as the green list, while $\delta$ determines the extent to which the selection probability of green tokens is amplified.
+We conduct an analysis on the HumanEval+ dataset by varying the values of $\delta$ and $\gamma$.
+Figure 3 shows the results in a bubble chart, revealing a clear trade-off—improving one metric often reduces the other.
+The circular markers represent the pass@1 and detection performance (AUROC) of STONE, while the triangular markers indicate those of SWEET.
+Among the four regions divided by two black dashed lines, the upper-right quadrant corresponds to the area where code quality and detection performance are both high, indicating a desirable balance between utility and detectability.
+We observe that while only a single configuration of SWEET falls into this balanced region, multiple configurations of STONE—spanning various combinations of $\gamma$ and $\delta$—lie within it.
+This suggests that STONE achieves a more balanced trade-off between functionality and watermark detectability compared to SWEET.
+Figure 3: The trade-off between pass@1 (Y-axis) and detectability (X-axis).
+We analyze the impact of the green list ratio $\gamma$ (indicated by marker color) and the watermark strength $\delta$ (indicated by marker size).
+Circular and triangular markers represent STONE and SWEET, respectively.
+6.2 Robustness under Adversarial Attacks
+Although a watermark is embedded in the code, a malicious user may remove it by modifying the code.
+Therefore, the watermark must remain robust against attacks like code modifications.
+We assess the resilience of watermarks of STONE against code alterations by applying two types of attacks: (1) a code refactoring attack using a commercial service [^1] https://codepal.ai/code-refactor and (2) a code paraphrasing attack using ChatGPT (GPT-4o) [^2] Prompt: Please paraphrase the following code..
+We conduct this analysis on the HumanEval+ dataset.
+Figure 4 shows the classification performance (i.e., detectability) in distinguishing between watermarked LLM-generated code and human-written code after these attacks are applied to code watermarked by STONE and SWEET.
+The experimental results show that STONE is more robust than SWEET against both types of attacks.
+Refactoring reorganizes code while keeping syntax boundaries, so STONE preserves many watermarks inserted into non-syntax tokens.
+Paraphrasing changes variable names and expressions, which directly correspond to the embedding targets of STONE, and thus decreases detectability.
+SWEET shows the opposite tendency because refactoring alters its watermarking targets.
+STONE achieves higher robustness under both attacks because it embeds watermarks into a syntax-filtered token space that remains stable across structural and lexical changes.
+Figure 4: We evaluate the robustness of watermarking methods by applying two types of adversarial perturbations—code refactoring and paraphrasing—to watermarked code.
+Following these attacks, we assess the effectiveness of watermark detection.
+We compare the performance of watermarks embedded using SWEET and STONE to determine their resilience under both attack scenarios.
 7 Related Work
-Research on code watermarking remains limited due to the unique challenges posed by structural constraints in code, particularly in maintaining functional correctness, detectability, and naturalness.
-Unlike natural language, code must adhere to strict syntactic and semantic rules, making the development of effective watermarking techniques more complex.
-Lee et al.
+Post-hoc detection methods (Xu and Sheng, 2024; Shi et al., 2024) identify machine-generated code after generation using statistical cues such as perplexity or syntax differences.
+In contrast, watermarking embeds provenance signals during generation to enable direct identification.
+This work focuses on the latter approach, developing a syntax-aware watermarking method for code generation.
+Among watermarking approaches, Lee et al.
 (2024) proposed SWEET, a method that selectively embeds watermarks in high-entropy tokens.
@@
 Our approach ensures that essential structural elements, like keywords and operators, remain unchanged, thereby preserving functional correctness without additional computational overhead.
-ACW (Li et al., 2024) and SrcMarker (Yang et al., 2024) adopted watermarking techniques based on code transformations, applying semantic-preserving modifications and variable name substitutions.
-Although these methods maintained program functionality, the introduced changes often result in unnatural code patterns, making the watermark more susceptible to removal by adversaries.
+ACW (Li et al., 2024a) and SrcMarker (Yang et al., 2024) employed watermarking techniques based on code transformations, such as semantic-preserving edits and variable renaming.
+While these approaches preserved program functionality, the resulting modifications often introduce unnatural code patterns, which can increase the risk of watermark removal by adversaries.
 8 Conclusion
-We propose STONE, a syntax-aware watermarking method that preserves the original functionality of code.
-Unlike existing methods that apply watermarks without considering syntax tokens—often causing syntax errors or functional failures—STONE restricts watermarking to non-syntax tokens, maintaining code structure and enhancing detection reliability.
-Additionally, we introduce CWEM, a comprehensive evaluation metric that assesses correctness, detectability, and naturalness.
-While correctness and detectability are well-studied, naturalness has been largely overlooked.
-By incorporating naturalness into the evaluation, CWEM provides a more balanced assessment of watermarking techniques.
-Experimental results show that STONE consistently outperforms existing methods across three programming languages, achieving higher CWEM scores.
-These findings demonstrate that syntax-aware watermarking improves detection performance while preserving code functionality, making it more practical for real-world applications.
+We present STONE, a syntax-aware watermarking method that preserves functional correctness by excluding syntax-critical tokens during watermark insertion.
+We also propose STEM, a unified evaluation framework integrating correctness, detectability, and imperceptibility for balanced assessment.
+Experiments on Python, C++, and Java show that STONE outperforms baseline methods without degrading code quality and reduces watermark detection time compared with entropy-based approaches.
+Evaluation under the STEM framework demonstrates balanced performance across weight configurations, and further analysis confirms that STONE maintains stable and robust performance under code modifications.
 Limitations
-This study introduces the STONE method, which preserves key functional tokens in code to maintain correctness.
-However, the current approach is not specifically optimized to enhance the naturalness of code during watermark insertion.
-Experimental results indicate that the naturalness score is lower than those achieved with methods such as KGW and EWD.
-Future work should explore refined token selection strategies to address this aspect.
-In our study, CWEM does not incorporate a measure for evaluating the persistence of the watermark when modifications, refactoring, or token substitutions alter the code.
-Future investigations should develop quantitative measures to assess watermark robustness under these conditions.
-We conduct experiments using the Qwen2.5-Coder model across three programming languages: Python, C++, and Java.
-The evaluation does not extend to additional large language model architectures, training environments, or parameter scales.
-Future studies should broaden the evaluation to encompass other model architectures, training strategies, and parameter scales in order to develop the scope and limitations of the approach.
-Appendix A Syntax Elements Coverage in Watermarking Tokens of SWEET
-In Table 5, we analyzed the SWEET methodology impact on token selection for watermarking using two metrics: (1) the proportion of tokens exceeding the entropy threshold among all tokens and (2) the proportion of syntax tokens among watermarking-selected tokens.
-Using SWEET optimal settings ($\gamma$=0.25, $\delta$=3.0, entropy threshold = 0.9), we examined threshold levels of [0.7, 0.8, 0.9, 1.0, 1.1].
-The results show an inverse relationship between threshold levels and token selection.
-Lower thresholds increase overall token selection while incorporating more syntax tokens, whereas higher thresholds reduce syntax token inclusion.
-These findings validate their selective watermarking approach, as it effectively targets high-entropy regions while preserving syntactic structures critical to text integrity.
-However, even at higher entropy thresholds, a non-negligible proportion of syntax tokens is still selected for watermarking.
-For the HumanEval+ dataset specifically, syntax tokens constituted 12.6% of watermarking tokens, with delimiters at 49.29%, whitespaces at 38.17%, keywords at 9.44%, types at 3.17%, and operators at 2.78%.
-This finding demonstrates that even in optimal settings, syntax tokens are inevitably selected as watermarking tokens.
-This distribution indicates that delimiters make up the majority of syntax tokens selected for watermarking, which can directly impact code execution.
-Table 6 illustrates the specific syntax tokens comprising the 12.6% identified in the HumanEval+ dataset.
-Delimiters, such as colons (‘:’) and commas (‘,’), are crucial for structuring code, and their modification can introduce syntax errors (e.g., omitting a colon in ‘for i in range(10):’ causes a parsing failure).
-While keywords and operators are selected less frequently, their modification can significantly alter program logic, as seen in cases like replacing ‘True’ with ‘False’ or ‘+’ with ‘-’.
-Given these risks, refining the token selection process to minimize disruptions to code execution, such as selectively excluding critical syntax tokens, could improve the practicality of syntax-aware watermarking.
-Entropy Threshold MBPP+ HumanEval+ HumanEvalPack - C++ HumanEvalPack - Java
-Watermarked Syntax Watermarked Syntax Watermarked Syntax Watermarked Syntax
-0.7 26.03 12.82 35.28 12.84 30.10 12.99 26.60 11.98
-0.8 21.40 12.59 32.67 12.24 25.93 12.35 22.47 11.06
-0.9 19.64 11.39 28.98 12.60 22.65 11.79 18.52 11.68
-1.0 16.59 11.39 24.84 11.83 18.96 10.56 16.13 11.15
-1.1 14.94 10.30 23.22 11.81 15.71 09.82 14.39 10.91
-Table 5: Syntax elements coverage in watermarking tokens of SWEET.
-Watermarked is the proportion of tokens that exceeds the entropy threshold among all tokens (watermarking tokens).
-Syntax is the proportion of watermarking tokens that belong to syntax elements.
-Bolded results correspond to the optimal entropy threshold setting used in SWEET.
-Category Examples
-Keywords def, or, import, for, True, assert, return, is, pass, None, False, in, not, if, from
+While STONE establishes a robust methodology for functionality-preserving watermarking, it has several limitations.
+First, the watermarking capacity is based on the density of non-syntactic tokens.
+The conventional coding styles in our benchmark datasets provided ample opportunities for watermark insertion.
+This efficacy, however, can be compromised in unconventional code, such as obfuscated scripts or solutions for code golf, where such tokens are sparse.
+Under these conditions, the embedded watermark may prove too sparse for reliable detection.
+Second, although our evaluation demonstrates resilience to general attacks like paraphrasing and refactoring, a more sophisticated threat model considers an adversary with specific knowledge of the STONE algorithm.
+Such an adversary may directly attack non-syntactic tokens by systematically renaming all variables and removing comments.
+This class of attack poses a more significant threat to watermark integrity than general modifications.
+Developing countermeasures against such targeted, algorithm-aware attacks remains a critical avenue for future research.
+Appendix A Syntax Elements in Python, C++, and Java
+Table 7 categorize the syntactic elements into five groups: keywords, whitespace, types, delimiters, and operators.
+The selection of syntactic elements reflects the distinct characteristics of each programming language.
+Appendix B Performance Analysis of CodeIP
 […]
-Types int, set, str
-Delimiters ‘.’,:, {, ‘,’, ], [, (, ) }
-Operators *, %, =, -, **, &, ^, /, +, |
-Table 6: Examples of syntax tokens selected as watermarked tokens in SWEET
-Appendix B Comparison of Syntax Elements in Python, C++, and Java
-This section provides examples of the syntactic elements in Python, C++, and Java that were utilized in our experiments.
-Table 7 categorize these elements into five groups: keywords, operators, delimiters, whitespaces, and types.
-The selection of syntactic elements reflects the distinct characteristics of each programming language.
+[…]
+[…]
+[…]
+[…]
+Table 5: Evaluation of CodeIP on four datasets. Correctness: pass@1; Detectability: AUROC; Imperceptibility: PPL gap (negative means degradation).
+Table 5 reports the reproduced results for CodeIP, a grammar-guided multi-bit watermarking method that employs a type prediction model trained on CodeSearchNet for grammar-aware token selection during code generation.
+All experiments are based on the original implementation and evaluated on four benchmarks: MBPP+, HumanEval+, HumanEvalPack-C++, and HumanEvalPack-Java.
+CodeIP maintains high detectability, achieving AUROC values above 0.94 across all datasets.
+This result indicates that grammar-based watermark insertion enables reliable message extraction.
+However, the strong detection capability is accompanied by severe degradation in correctness and imperceptibility.
+The pass@1 scores decrease sharply, approaching zero on HumanEval+ and HumanEvalPack-C++, which shows that syntactic validity enforced by the type predictor does not ensure functional accuracy.
+The large perplexity gap of up to +4.9 indicates that the interaction between watermarking and type-prediction constraints distorts the token distribution, causing deviations from the model’s natural generation pattern.
+Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ HEP-C++ HEP-Java
+$(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.270 0.296 0.122 0.018
+(0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.241 0.274 0.095 0.041
+(0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.381 0.361 0.267 0.145
+(0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.325 0.331 0.174 0.084
+Table 6: STEM scores of CodeIP under different $(\alpha,\beta,\zeta)$ on MBPP+, HumanEval+, HumanEvalPack-C++, and HumanEvalPack-Java.
+Table 6 presents the STEM scores of CodeIP under different weight settings.
+Across all configurations, detectability remains consistently high, while correctness- or imperceptibility-focused STEM scores are significantly lower than those of other methods.
+Even under weight settings emphasizing correctness or imperceptibility, CodeIP does not achieve balanced performance, showing limited robustness under varying evaluation preferences.
+These findings demonstrate that maximizing watermark detectability without syntax awareness severely degrades code quality.
+In contrast, STONE embeds watermarks only within non-syntax tokens and achieves stable performance across correctness, detectability, and imperceptibility as captured by the STEM metric.
 Category Python C++ Java
@@
 […]
-Table 7: Comparison of five types of syntax elements across Python, C++, and Java.
-Appendix C Trade-off Between Correct and Detect
-We analyze the trade-off between pass@1 and AUROC based on watermarking strength by adjusting two parameters: $\gamma$, which controls the ratio between the green list and red list, and $\delta$, which determines how often the model selects tokens from the green list.
-Figure 3 shows the results in a bubble chart, revealing a clear trade-off—improving one metric often reduces the other.
-The bottom-right section of each subfigure highlights larger light-blue bubbles, where high delta and low gamma values produce higher AUROC but lower pass@1.
-A smaller gamma shrinks the green list, limiting the token options available during generation.
-At the same time, a higher delta increases the likelihood of selecting tokens from the green list.
-This combination strengthens the watermark signal and raises AUROC but generates less natural code, reducing pass@1.
-Achieving the right balance between code quality and watermark detectability requires careful tuning of gamma and delta.
-Maintaining code quality and correctness remains crucial when embedding watermarks, so relying on both a low gamma and a high delta together creates undesirable results.
-The areas shaded in green highlight points where pass@1 and AUROC achieve a proper balance.
-The double circles outlined in black represent the performance reported in Table 2.
-For the MBPP+ dataset, these points correspond to $\delta$ = 1.0 and $\gamma$ = 0.5, while for the HumanEval+ dataset, they correspond to $\delta$ = 0.5 and $\gamma$ = 0.5.
-The trends in the green-shaded region suggest that using a lower delta value generally results in a better balance between pass@1 and AUROC performance.
-Figure 3: The tradeoff between Correct and Detect.
-We analyze the impact of the green list ratio $\gamma$ (indicated by marker color) and the logit adjustment factor $\delta$ (indicated by marker size).
-Each point represents the functional correctness of code watermarked using the STONE (Y-axis) and the performance of classifying it against human-written code (X-axis).
-The green-shaded area highlights points where Correct and Detect are well balanced.
-Double circles with an outer black ring mark the performance points reported in Table 2.
+Table 7: Comparison of five categories of syntax elements across Python, C++, and Java.
+Appendix C Experimental Results on Llama-3.1-8B
+Tables 8 and 9 summarize the experimental results obtained with the Llama-3.1-8B model on MBPP+ and HumanEval+.
+STONE achieves the highest STEM scores across all datasets and weight settings, demonstrating that the non-syntax watermarking strategy generalizes well to larger models.
+EWD shows slightly higher detectability on HumanEval+, which appears to result from an interaction between its uniform insertion strategy and the sharper probability distributions of Llama-3.1.
+In this setting, the cumulative bias introduced by EWD is more effectively reinforced across long and diverse sequences, leading to a small gain in detectability.
+Such an effect is not observed under Qwen2.5-Coder-7B, whose flatter distributions yield comparable results between EWD and SWEET.
+Overall, STONE preserves stronger correctness and achieves a more balanced trade-off, resulting in the highest overall STEM scores.
+Across different weight configurations, STONE maintains stable performance even when correctness or imperceptibility is prioritized.
+The method remains competitive under detectability-oriented settings without functional degradation.
+This consistency confirms that syntax-aware watermarking effectively preserves code quality while sustaining robust watermark signals.
+Method MBPP+ HumanEval+ Correct.
+Detect.
+Imperc.
+Correct.
+Detect.
+Imperc.
+KGW 0.414 0.831 0.953 0.323 0.507 0.989 EWD 0.414 0.931 0.953 0.323 0.755 0.989 SWEET 0.404 0.927 0.968 0.354 0.711 0.978 STONE 0.446 0.945 0.954 0.372 0.741 0.983
+Table 8: Experimental results on MBPP+ and HumanEval+ (Llama-3.1-8B): Correctness, Detectability, and Imperceptibility.
+  Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ KGW EWD SWEET STONE KGW EWD SWEET STONE $(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.733 0.766 0.766 0.782 0.606 0.689 0.681 0.699 (0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.653 0.678 0.676 0.698 0.535 0.598 0.599 0.617 (0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.757 0.807 0.806 0.823 0.582 0.706 0.688 0.709 (0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.788 0.813 0.817 0.825 0.702 0.764 0.755 0.770
+Table 9: STEM scores on MBPP+ and HumanEval+ (Llama-3.1-8B) under different weight settings $(\alpha,\beta,\zeta)$.
+The left column lists weights for correctness, detectability, and imperceptibility.
+Bold indicates the highest score for each weight configuration.
+Entropy Threshold MBPP+ HumanEval+ HEP-C++ HEP-Java
+Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%)
+0.7 26.03 12.82 35.28 12.84 30.10 12.99 26.60 11.98
+0.8 21.40 12.59 32.67 12.24 25.93 12.35 22.47 11.06
+0.9 19.64 11.39 28.98 12.60 22.65 11.79 18.52 11.68
+1.0 16.59 11.39 24.84 11.83 18.96 10.56 16.13 11.15
+1.1 14.94 10.30 23.22 11.81 15.71 9.82 14.39 10.91
+Table 10: Token selection statistics of SWEET across four datasets.
+Selected Tokens (%) denotes the proportion of tokens exceeding the entropy threshold out of all tokens, and Syntax Tokens (%) represents the percentage of syntax-related tokens among those selected.
+Bold values indicate the optimal entropy threshold used in SWEET.
+  Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ HEP-C++ HEP-Java KGW EWD SWEET STONE KGW EWD SWEET STONE KGW EWD SWEET STONE KGW EWD SWEET STONE $(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.775 0.819 0.787 0.848 0.694 0.763 0.754 0.781 0.730 0.750 0.735 0.780 0.642 0.675 0.631 0.715 (0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.706 0.739 0.716 0.778 0.664 0.716 0.709 0.732 0.692 0.706 0.697 0.741 0.578 0.603 0.577 0.648 (0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.789 0.856 0.807 0.881 0.651 0.755 0.743 0.780 0.703 0.733 0.711 0.768 0.618 0.668 0.618 0.716 (0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.830 0.863 0.838 0.883 0.767 0.810 0.808 0.830 0.796 0.811 0.796 0.833 0.730 0.755 0.699 0.781
+Table 11: STEM scores under different weight settings $(\alpha,\beta,\zeta)$.
+The left column lists weights for correctness, detectability, and imperceptibility.
+Bold indicates the best score in each row.
+Appendix D Comprehensive Evaluation across STEM Weight Configurations
+Table 11 reports representative STEM variants using equal weights and three dimension-focused settings.
+We further explore all valid weight configurations where $\alpha,\beta,\zeta\in\{0.0,0.1,\dots,1.0\}$ and $\alpha+\beta+\zeta=1$, resulting in $66$ combinations.
+For each configuration, we compute the weighted STEM score and record the method with the highest value.
+Across all datasets, STONE performs consistently well throughout the entire weight space, ranking first in $97.0\%$ of configurations on MBPP+, $90.9\%$ on HumanEval+, $98.5\%$ on HumanEvalPack-C++, and $95.5\%$ on HumanEvalPack-Java.
+This consistency indicates that the advantage of STONE does not depend on specific weight choices but generalizes across diverse evaluation preferences, demonstrating robust performance under varying priorities of correctness, detectability, and imperceptibility.
+Appendix E Syntax Token Coverage in SWEET Watermarking
+Table 10 presents an analysis of token selection by SWEET using two measures: (1) the proportion of selected tokens exceeding the entropy threshold among all generated tokens, and (2) the proportion of syntax-related tokens among those selected.
+We use the optimal settings ($\gamma$ = 0.25, $\delta$ = 3.0, entropy threshold = 0.9)) and evaluate threshold values in the range [0.7, 0.8, 0.9, 1.0, 1.1].
+The results show an inverse relationship between the threshold and token coverage: lower thresholds select more tokens overall but include more syntax elements, whereas higher thresholds reduce syntax token inclusion.
+This pattern shows that SWEET effectively targets high-entropy regions but does not completely avoid syntax-related tokens.
+Even under the optimal configuration, a notable portion of syntax tokens remains among those selected for watermarking.
+In the HumanEval+ dataset, syntax tokens account for 12.6% of all selected tokens, including delimiters (49.29%), whitespace (38.17%), keywords (9.44%), types (3.17%), and operators (2.78%).
+This indicates that delimiters and whitespace dominate syntax-related selections, which directly affects code execution.
+Category Examples
+Keywords def, or, import, for, True, assert, return, is, pass, None, False, in, not, if, from
+[…]
+Types int, set, str
+Delimiters . , : , {, }, [, ], (, )
+Operators *, %, =, -, **, &, ^, /, +, |
+Table 12: Examples of syntax tokens selected for watermarking in SWEET.
+Table 12 lists representative examples.
+Delimiters such as colons (‘:’) and commas (‘,’), essential for code structure, cause parsing errors when modified (e.g., omitting a colon in ‘for i in range(10):’ leads to a failure).
+Although keywords and operators appear less frequently, their modification can still alter program behavior, for example, replacing ‘True’ with ‘False’ or ‘+’ with ‘-’.
+Minimizing the inclusion of such syntax-critical tokens can make syntax-aware watermarking more reliable and less disruptive to functionality.
+Appendix F Comparison with Non-watermarked Code
+Method MBPP+ HumanEval+
+No watermark 0.571 0.595
+KGW 0.499 0.573
+EWD 0.499 0.573
+SWEET 0.502 0.574
+STONE 0.571 0.587
+Table 13: Correctness comparison between non-watermarked and watermarked code. Correctness is measured as pass@1 on MBPP+ and HumanEval+.
+We evaluate the functional impact of watermarking by comparing the correctness of code generated with and without watermark insertion.
+Non-watermarked code defines the upper bound of functional performance.
+STONE achieves scores nearly identical to this bound, showing that watermark insertion has minimal influence on code execution.
+Among all watermarking methods, STONE reduces correctness the least, which indicates that it embeds provenance information effectively while preserving the original program behavior.